OpenStudy (anonymous):

Find the value of the variable(s). If your answer is an integer, leave it in simplest radical form. The triangle is not drawn to scale.

OpenStudy (anonymous):

\[a) x=17, y=34\sqrt{3}\] \[b) x=34, y=17\sqrt{3}\] \[c) x=34\sqrt{3}, y=17\] \[d) x=17\sqrt{3},y=34\]